40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Wed Apr 28 10:16:09 2010, Frank, Lab Infrastructure, General, lab cleanliness - latest particle counts AdhikariLab_data_repair.png
    Reply  Wed Apr 28 11:19:36 2010, Alastair, Lab Infrastructure, General, lab cleanliness - latest particle counts 
    Reply  Wed Apr 28 13:21:11 2010, Alastair, Lab Infrastructure, General, lab cleanliness - latest particle counts 
    Reply  Wed Apr 28 14:57:14 2010, rana, Lab Infrastructure, General, lab cleanliness - latest particle counts 
    Reply  Wed Apr 28 19:10:57 2010, Dmass, Lab Infrastructure, General, lab cleanliness - latest particle counts 
Message ID: 738     Entry time: Wed Apr 28 11:19:36 2010     In reply to: 737
Author: Alastair 
Type: Lab Infrastructure 
Category: General 
Subject: lab cleanliness - latest particle counts 

Quote:

really bad 

...

they tried to fix the leaks and fixed a lot of small ones but unfortunately the didn't fix the large ones. I checked five of the large ones and four of them still exist...
The general cleanliness is very bad, especially since a couple of month. You don't have to be an expert to see how bad it is. For some people it seems to be normal to step over the black (non-sticky anymore) mats and not to remove the (obviously) dirty and unfunctional top layer. And keeping the HEPAs on the 35W table off since a week(!) (or longer?)  doesn't help much to keep this very sensitive area clean...

Hey Frank, do we know when they're planning on being done down here?  The guy I spoke to said that they'd re-check for leaks once they turned the air back on.

Definitely needs cleaned, but I was hoping they would be done this week and we could clean after they had gone.  If it's going to take longer than that maybe we should arrange an intermediate cleaning.

There are some guys in the lab this morning switching out the smoke detectors for the fire system so I'll let them into the labs they need in.  They're turning off the air for five minutes to do this but I assume this isn't a problem since it's been on and off this week already.

 Does anyone know why the spectrum analyser is on it's side on the floor?

ELOG V3.1.3-